### v1

# NeuSemSlice: Towards Effective DNN Model Maintenance via Neuron-level Semantic Slicing

Abstract.
Deep Neural networks (DNNs), extensively applied across diverse disciplines, are characterized by their integrated and monolithic architectures, setting them apart from conventional software systems. This architectural difference introduces particular challenges to maintenance tasks, such as model restructuring (e.g., model compression), re-adaptation (e.g., fitting new samples), and incremental development (e.g., continual knowledge accumulation). Prior research addresses these challenges by identifying task-critical neuron layers, and dividing neural networks into semantically-similar sequential modules. However, such layer-level approaches fail to precisely identify and manipulate neuron-level semantic components, restricting their applicability to finer-grained model maintenance tasks.
In this work, we implement NeuSemSlice, a novel framework that introduces the semantic slicing technique to effectively identify critical neuron-level semantic components in DNN models for semantic-aware model maintenance tasks. Specifically, semantic slicing identifies, categorizes and merges critical neurons across different categories and layers according to their semantic similarity, enabling their flexibility and effectiveness in the subsequent tasks.
For semantic-aware model maintenance tasks, we provide a series of novel strategies based on semantic slicing to enhance NeuSemSlice. They include semantic components (i.e., critical neurons) preservation for model restructuring, critical neuron tuning for model re-adaptation, and non-critical neuron training for model incremental development. A thorough evaluation has demonstrated that NeuSemSlice significantly outperforms baselines in all three tasks.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
1. Introduction
In the evolving landscape of artificial intelligence, DNNs represent a dynamic shift towards data-driven programming paradigms, extensively adopted across various fields (Kollias et al., 2018; Bansal et al., 2018; Théate and Ernst, 2021). Analogous to how traditional software evolves to meet shifting requirements and enhance functionality, effective maintenance of DNNs is crucial for combating model drift and ensuring sustained accuracy and reliability as they are continuously exposed to new and changing data. Unlike traditional software systems characterized by their modular and discrete structures, DNNs are typically designed as cohesive, monolithic entities. The unique architecture introduces distinct and challenging maintenance tasks, including model restructuring, re-adaptation, and incremental development. They are pivotal elements for the operational robustness and longevity of neural network-based systems in a rapidly advancing technological milieu (AI model maintenance: ensuring accuracy, scalability, and reliability, 2023; AI & ML Model Maintenance, 2023; AI Maintenance: A Robustness Perspective, 2023).
Despite various proposals in the literature for maintaining DNN models like the DeepArc framework (Ren et al., 2023), these methods continue to face several limitations. One major challenge with traditional strategies, which depend largely on periodical model retraining, is their ineffectiveness in combating the unavoidable problem of model drift. DeepArc addresses this problem by modularizing neural networks, organizing layers into modules based on their semantic similarities. This strategy streamlines maintenance tasks, facilitating the pruning and fine-tuning of specific modules or layers for more efficient updates. However, this approach is constrained by its focus on layer-level manipulation. A more refined strategy that decomposes semantic components at a finer level, such as individual neurons, could enhance model restructuring and adaptation. Specifically, selectively pruning and tuning a small number of neurons might offer better effectiveness. Furthermore, the incremental development of models (Wang et al., 2024) requires preserving the semantics and functions of previous versions while introducing new features. Given that semantics and functions in DNNs are primarily structured around neurons (Du et al., 2019; Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021; Shrikumar et al., 2017), the method of identifying and updating whole layers falls short of preserving the original semantics or functions during the incremental development process.
Acknowledging the existing limitations, our goal is to develop techniques that enhance DNN model maintenance. Drawing inspiration from the significant benefits of dynamic slicing in the maintenance and evolution of traditional software (Beszédes et al., 2001; Jiang et al., 2014), we adopt similar slicing techniques for DNNs (Zhang et al., 2020a; Li et al., 2023; Xie et al., 2022; Sun et al., 2022; Liu et al., 2018). Specifically, dynamic slicing excels in identifying and isolating essential code segments for implementing changes without causing unintended effects. By analogy, DNN slicing provides a systematic way to identify and separate semantic components at the neuron level. This approach is poised to greatly enhance DNN model maintenance, allowing for precise modifications and improvements with a focus on finer detail.
Our work. We introduce the NeuSemSlice framework, building on the concept of semantic decomposition through DNN slicing, to innovate model maintenance practices. This framework features two main components: semantic slicing and semantic-aware model maintenance. The former facilitates the precise identification of semantic elements within DNNs, such as critical neurons, at a granular level. Following this, the latter delivers tailored strategies for model maintenance tasks that demand an in-depth semantic comprehension, enabling effective and targeted strategies.
The semantic slicing module is designed to accurately identify critical neurons across the entire search space by employing a category-specific [^1] To be precise, we identify critical neurons within the network model that have a significant impact on the output of each output neuron, where different categories refer to different output neurons. For example, in classification tasks, categories can be determined based on the labels in the dataset; in regression tasks, it can be considered a single-category task; in other types of tasks, different output neurons are treated as different categories. strategy for optimal performance. Specifically, we compute neuron contribution scores (Pei et al., 2017; Ish-Horowicz et al., 2019; Xie et al., 2022; Li et al., 2023; Molchanov et al., 2019) for each neuron category individually, selecting a specific range of critical neurons for each category before combining them across the entire dataset. Given the potential variability in the optimal range of critical neurons across different categories and layers—which may hinder effective combination—we refine this process to focus on neurons that encapsulate the layer’s full semantic range. This is achieved by assessing the similarity between the characteristics of these chosen neurons and the entire layer. The subsequent merging strategy further enhances the utility and relevance of the identified critical neurons.
The semantic-aware model maintenance module, utilizing semantic slicing, provides the following maintenance tasks for DNN models: ❶ Model Restructuring, where only critical neurons are retained to balance predictive precision with reduced storage needs. ❷ Model Re-adaptation, focusing on training critical neurons to address prediction errors, thus simplifying the process by tuning fewer parameters. ❸ Model Incremental Development, which stabilizes critical neuron parameters to prevent catastrophic forgetting during the integration of new knowledge, facilitating incremental development without extensive reliance on the original dataset.
We conduct extensive experiments to evaluate the performance of NeuSemSlice. For model restructuring, NeuSemSlice enhances the efficiency of existing model compression techniques by an average of 13.46%, while also improving compression rate. For model re-adaptation, NeuSemSlice only trains an average of 61.49% of the parameters, enhancing both accuracy and efficiency. For incremental model development, NeuSemSlice surpasses all other methods, boosting average accuracy by over 40% compared to baseline.
Contributions. The contribution of this paper is summarized as follows: ❶ We propose NeuSemSlice, including semantic slicing for finer-grained neuron-level decomposition, and semantic-aware model maintenance, offering novel and effective strategies for semantic-aware DNN model maintenance tasks. ❷ We detail a sophisticated semantic slicing technique that employs a category-specific strategy to accurately identify and combine critical neurons across the dataset, refining the process of neuron selection based on semantic range to optimize model performance. ❸ We tackle the challenging maintenance tasks via the semantic-aware model maintenance module, which significantly advances model restructuring, adaptation, and incremental development tasks. ❹ We conduct extensive experiments on three model maintenance tasks, showcasing the superior performance of NeuSemSlice.
2. Background and Related Work
2.1. Deep Neural Networks
A DNN typically consists of multiple layers of neurons (Tishby and Zaslavsky, 2015), which can be formally defined as follows.
Definition 0.
A Deep Neural Network (DNN) $f$ consists of $L$ multiple layers $\left\langle l_{0},l_{1},\ldots,l_{L-1}\right\rangle$, where $l_{0}$ is the input layer, $l_{L-1}$ is the output layer, and $l_{1},\ldots,l_{L-2}$ are hidden layers. The inputs of each layer are the outputs of the previous layer.
In this work, we mainly focus on the classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}$ is a set of inputs and $\mathcal{Y}$ is a set of classes. Given an input $x\in\mathcal{X}$, we use $f^{l}(x)$ to represent the internal features extracted by the layer $l$ and $f_{n}(x)$ to represent the internal features extracted by the neuron $n$ (i.e., the output values of neurons at $l$). $\hat{y}$ represents the DNNs output (i.e., $f(x)$) and $y$ is the label of the given input $x$.
There usually exists redundancy in DNNs which refers to components within the network that do not significantly contribute to its performance (Cheng et al., 2015; Xie et al., 2022). This redundancy can manifest in various forms, such as redundant neurons and redundant layers. Nevertheless, determining which components are redundant and can be removed without impacting performance requires sophisticated analysis and careful validation.
2.2. DNN Model Maintenance
This concept, akin to the established practices of traditional software maintenance, addresses the evolving needs within the software engineering community for more efficient and adaptable DNN systems. DeepArc (Ren et al., 2023) has introduced a new neural network architecture strategy to cut model maintenance costs. This approach identifies semantically similar layers as redundant, allowing for model restructuring by pruning these layers. Only non-redundant layers are fine-tuned on the target data, based on calculated semantic similarities of layer representations: Given a dataset $D$, the matrix $M^{l_{i}}$ is denoted as the similarity matrix for the $i$-th network layer $l_{i}$, where
(1) $M^{l_{i}}[a,b]=f^{l_{i}}(x_{a})\cdot f^{l_{i}}(x_{b}),$
with $x_{a}$ and $x_{b}$ representing any two inputs from the dataset $D$. Intuitively, $M^{l_{i}}$ indicates the embedding landscape of the dataset on the $i$-th layer and the shape of the matrix is solely dependent on the size of the dataset. Similar embedding landscapes (i.e., similar matrix) in two layers indicate they extract comparable semantic information from inputs. With the matrix $M^{l}$ for each layer $l$, DeepArc then follows the centered kernel alignment (Kornblith et al., 2019) (CKA) metric to calculate the similarity between $M^{l_{i}}$ and $M^{l_{j}}$ to represent the semantic similarity between layer $f^{l_{i}}$ and $f^{l_{j}}$.
(2) $\text{sim}_{\text{sem}}(f^{l_{i}},f^{l_{j}})=\text{CKA}(M^{l_{i}},M^{l_{j}})$ $=\frac{M^{l_{i}}\cdot M^{l_{j}}}{||M^{l_{i}}||\cdot||M^{l_{j}}||}.$
A higher $\text{sim}_{\text{sem}}$ indicates greater semantic similarity between consecutive layers, indicating redundancy. Removing these redundant layers aids in model restructuring, while focusing retraining efforts on only the essential, non-redundant layers facilitates model re-adaptation. This strategy, however, only targets redundant layers without considering redundant neurons, causing limitations to be detailed in Section 2.4.
2.3. DNNs Slicing
DNNs, essentially programs made up of artificial neurons, have seen recent innovations in “slicing” techniques. These methods identify key neurons within DNNs that are crucial for various tasks (Xie et al., 2022; Zhang et al., 2020a). The essence of DNN slicing research revolves around using a neuron contribution metric to evaluate the impact of each neuron on a dataset $D$. A neuron selection strategy is then applied to pick out critical neurons based on their contribution to the network’s functionality.
Specifically, the process could be formulated as follows: we calculate $\mathbf{C}_{n}^{l,D}$ for each neuron $n$ in layer $l$ based on a dataset $D$, the neuron contribution of the layer $l$ composed of neuron set $\mathbf{N}^{l}$ could be represented as $\mathbf{C}^{l,D}=\{\mathbf{C}_{n_{0}}^{l,D},\mathbf{C}_{n_{1}}^{l,D},...,\mathbf{C}_{n_{|\mathbf{N}^{l}|-1}}^{l,D}\}$. We further determine the selection of critical neurons according to the calculated neuron contribution and denote the selected critical neurons as $\mathbf{\Omega}^{l,D}_{k}$:
(3) $\mathbf{\Omega}_{k}^{l,D}=\text{top}\-k(\mathbf{N}^{l,D},\mathbf{C}^{l,D})$
where $\text{top}\-k(\cdot)$ represents the top $k$ maximum instances of the input set $\mathbf{N}^{l}$ based on certain metrics. We can further depict the critical neurons of the model $f$ as $\mathbf{\Omega}_{k}^{D}=\{\mathbf{\Omega}_{k}^{l_{0},D},\mathbf{\Omega}_{k}^{l_{1},D},...,\mathbf{\Omega}_{k}^{l_{L-1},D}\}$. Note the proportion of selected critical neurons is directly controlled by hyperparameter $k$.
The predominant approaches of neuron contribution metric design for neuron $n$ are as follows: ❶ Avg (Average of Neuron Activation (Pei et al., 2017)). We compute the mean activation value for each neuron across the dataset, defined as: $avg_{n}(D)=\frac{\sum_{x\in D}f_{n}(x)}{|D|}.$Var (Variance of Neuron Activation (Ish-Horowicz et al., 2019)). We compute the activation value variance for each neuron using the formula: $var_{n}(D)=\frac{\sum_{x\in D}(f_{n}(x)-\overline{f_{n}(x)})^{2}}{|D|}.$Ens (Ensemble of Neuron Activation). With $avg_{n}(D)$ and variance $var_{n}(D)$ of each neuron in the neural network concerning dataset $D$, we normalize the two values to [0,1]. ❹ DeepLIFT (Interpretability-based Score (Xie et al., 2022; Li et al., 2023)). Given its interpretability, our analysis predominantly utilizes DeepLIFT (Shrikumar et al., 2017) for computations. In specific, DeepLIFT assigns contribution scores to each neuron based on the gradient during the backward propagation process. ❺ Taylor (Pruning-oriented Score (Molchanov et al., 2019)). Taylor utilizes model weights and gradient information to assess the contribution of neurons.
Leveraging the current neuron contribution metrics, there are intuitively two neuron selection strategies. The global neuron selection method calculates neuron contribution scores across the entire dataset $D$, followed by a selection of a specific range of neurons. In contrast, the category-wise neuron selection approach computes scores for each category of data ($D_{c_{0}}$, $D_{c_{1}}$, $\ldots$) separately, identifies a range of critical neurons for each category and layer (i.e., identifying $\mathbf{\Omega}_{k}^{l,D_{c_{0}}},\mathbf{\Omega}_{k}^{l,D_{c_{1}}},\ldots$), and ultimately merges these category-wise critical neurons for the entire dataset as a whole ($\mathbf{\Omega}_{k}^{l,D}$).
Figure 1. Overview of NeuSemSlice Framework. In semantic slicing, neuron contributions are highlighted in $purple$, with darker shades signifying greater contributions. Critical neurons are indicated in $orange$. For semantic-aware model maintenance, we preserve only these critical neurons. During model re-adaptation, we train solely the critical neurons, marked in $green$. For incremental development, only the non-critical neurons are trained, shown in $blue$.
2.4. Motivation
The state-of-the-art technique exhibits three noticeable limitations in addressing the challenging tasks of DNN model maintenance: ❶ Lack of granularity. Existing technique is confined to layer level, which lacks the flexibility for optimal model restructuring and adaptation on the finer neuron level. ❷ Ineffectiveness for smaller models. The existing technique is ineffective for smaller DNNs, crucial for on-device deployments, due to their lack of layers with similar semantics. This limits the ability to downsize by removing redundant layers for restructuring or adaptation. ❸ Absence of incremental model development. This crucial aspect of model maintenance is not adequately addressed. The current technique is unable to identify key components at the neuron level (Du et al., 2019), which is essential for the model’s functions, leading to challenges in developing new functions while preserving existing ones with only layer-level adjustments.
2.5. NeuSemSlice Overview
To address the issues, we propose NeuSemSlice which focuses on neuron-level DNNs semantic decomposition and involves slicing DNNs to identify and isolate critical neuron components, facilitating a variety of model maintenance tasks. Figure 1 shows a detailed workflow of our NeuSemSlice framework, which mainly consists of two modules: Semantic Slicing and Semantic-aware Model Maintenance.
Semantic slicing. This module initially identifies and extracts crucial neurons across the entire model. These neurons are representative of the semantics inherent in each layer of the model. By focusing on a more detailed level, it overcomes the initial two limitations, enabling more precise restructuring and adaptation. Additionally, it improves the ability to identify redundant neurons in the smaller layers of models.
Semantic-aware model maintenance. By utilizing the aggregated critical neurons in the model, this approach supports various applications, offering more robust solutions for long-term DNN model maintenance. Beyond restructuring and re-adapting the model, this method also addresses incremental model development for the first time. It specifically counteracts catastrophic forgetting, which occurs when an AI model learns new information, by stabilizing the critical neurons’ parameters. This enables the model to incrementally develop functions with limited access to the original training data.
In Section 3, we will provide a detailed introduction to the design methodology behind Semantic Slicing. Following that, in Section 4, we will discuss Semantic-aware Model Maintenance: the application of Semantic Slicing in the context of model maintenance tasks.
3. Semantic Slicing
As discussed in Section 2.3, various neuron contribution metrics and two possible neuron selection strategies can be adopted for neuron-level DNNs slicing. Yet, how well these methods perform in model maintenance tasks has not been fully examined. To address this, Section 3.1 details a pilot study designed to evaluate their effectiveness and efficiency in maintaining models. Following these experiments, we present our approach to semantic slicing in Section 3.3, guided by the insights gained.
3.1. A Pilot Study on Existing DNNs Slicing Strategies
Global Slicing. One possible strategy is the global approach, which calculates contribution metrics across the entire dataset and then selects the top 10% (i.e., set $k$ in Equation 3 as 10%) of neurons, as indicated by these metrics, as critical neurons. This process, outlined in Section 2.3, results in five distinct sets of critical neurons identified using different metrics. We then devise a strategy to evaluate the importance of each set to determine the most effective contribution metric. Specifically, to assess and compare the significance of the critical neurons identified by various metrics, we adopt the method from Xie et al. (Xie et al., 2022) by zeroing the outputs of these neurons and observing the impact on model predictions. Greater changes in predictions suggest the higher neuron significance.
The study is conducted on CIFAR-10 (Krizhevsky et al., 2009) and training VGG-16 (Simonyan and Zisserman, 2014) for classifying. Table 1 shows that in a global setting, all neuron contribution metrics have limitations. Surprisingly, masking neurons deemed less critical often leads to more significant accuracy losses, suggesting a misalignment between globally calculated contributions and the neurons’ true importance. For example, using the Avg metric, masking 20$\rightarrow$30of neurons has a greater impact than masking the top 10$\rightarrow$20%, as highlighted by the green arrows in the table. The results suggest that global slicing is not effective for model maintenance tasks.
Table 1. Model accuracy after masking selected critical neurons globally calculated by different contribution metrics. The anticipated trend in model accuracy is an increase (signified by $\uparrow$), correlating with the diminishing contribution of masked neurons. Nonetheless, within the context of global slicing, numerous anomalous instances of decreased model accuracy are observed, as indicated by $\downarrow$.
VGG16-CIFAR10
performance
after masking
[…]
0%->10% 42.85% 46.78% 50.64% 25.78% 41.29%
10%->20% 65.45% 76.91% 62.27% 36.83% 47.94%
[…]
Table 2. The model accuracy averaged by category after masking critical neurons derived by various metrics in categorical settings. The anticipated trend in model accuracy is an increase (signified by $\uparrow$), correlating with the diminishing contribution of masked neurons.
VGG16-CIFAR10
performance
after masking
[…]
0%->10% 0.00% 0.09% 0.00% 0.30% 11.80%
10%->20% 0.07% 5.16% 0.12% 8.48% 27.94%
20%->30% 1.25% 11.02% 2.49% 21.49% 48.24%
30%->40% 31.39% 37.81% 19.58% 59.12% 53.90%
40%->50% 55.45% 58.63% 57.61% 80.79% 62.64%
Category-wise Slicing. Another possible strategy is the category-wise approach. Acknowledging that neuron contributions can vary widely between categories, which may limit the effectiveness of global DNN slicing, we conduct a category-specific analysis. We calculate the five metrics for each category and then conduct masking experiments on data from each category (i.e., zeroing the outputs of neurons deemed critical for that category) to assess accuracy on a per-category basis. The averaged and individual category results are shown in Tables 2 and 3, respectively. Masking the top 10% of neurons shows that certain metrics (e.g., Avg and Ens) drastically lowered classification performance, nearly to zero. This demonstrates that a category-specific approach is more effective in identifying critical neurons than that of the global-based apporach.
3.2. Identified Challenges for Semantic Slicing
Identifying critical neurons for model maintenance involves two main challenges: the optimal proportion of critical neurons varies by category and by layer. ❶ Different category-wise distribution. Integrating category-specific critical neurons into a set for the entire dataset is complex as the ideal proportion of these neurons differs across categories. For example, masking 10% to 20% of critical neurons in Category 4 hardly affects performance, suggesting redundancy. In comparison, masking 30% to 40% in Category 1 significantly impacts performance, indicating their importance for that category. ❷ Different layer-wise distribution. The distribution of critical neurons changes across layers, often becoming sparser in higher layers. Current methods propose a uniform selection range for all categories and layers, which can either overlook crucial neurons or include too many non-critical ones, affecting accuracy and efficiency in model maintenance. Therefore, determining the optimal range of critical neurons for each category and layer is essential for effective DNN slicing.
However, achieving the goal is a challenging endeavor. The challenge lies in the impracticality of conducting an exhaustive search across all possible combinations of critical neuron proportions for each layer and category to assess performance after merging. Both the exhaustive search and the post-merging evaluation are inefficient in their current forms. To address this, we propose Semantic Slicing, to be detailed next in Section 3.3.
Table 3. Model performance on selected categories after masking selected critical neurons calculated by different contribution metrics. The anticipated trend in model accuracy is an increase (signified by $\uparrow$), correlating with the diminishing contribution of masked neurons.
VGG16-CIFAR10 (DeepLIFT)
[…]
0%->10% 0.00% 3.00% 0.00% 0.00%
10%->20% 0.00% 81.80% 0.00% 0.00%
20%->30% 13.30% 80.40% 31.70% 5.90%
30%->40% 33.20% 90.20% 39.20% 86.70%
40%->50% 78.50% 95.80% 90.80% 96.30%
3.3. The Design of Semantic Slicing
The essence of existing DNNs slicing methods revolves around employing a contribution metric and then taking a neuron selection strategy to determine the range of critical neurons. Compared with these approaches, the key novelty of semantic slicing lies in the semantic-based neuron selection strategy. The aim of semantic slicing is to address the issue of determining the appropriate $k$ value in Equation 3 for various layers and categories. To address this issue, we here recognize that the crux of choosing critical neurons lies in selecting those neurons whose semantics are adequately representative of the entire layer.
Problem Definition. Inspired by prior layer similarity evaluation works (Nguyen et al., 2020; Ren et al., 2023), we formally reformulate the task of selecting top $k$ neurons for each layer and category as follows. For each layer $l$ and category $c$, identify $k$ neurons such that they fulfill the condition: $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$, where $\Theta$ represents a unified threshold applicable across various categories and layers, and $\Omega_{k}^{l,c}$ is the selected neuron set. Employing the CKA metric as in Equation 2, we calculate a similarity metric in the range of [0, 1] as follows:
(4) $min|\mathbf{\Omega}^{l,c}|\quad\text{subject to}\quad\frac{M^{\mathbf{\Omega}^{l,c}}\cdot M^{l,c}}{||M^{\mathbf{\Omega}^{l,c}}||\cdot||M^{l,c}||}>\Theta,\vspace{-0.4cm}$
where $\mathbf{\Omega}^{l,c}$ is a neurons subset of layer $l$ selected on category $c$, and $M^{l,c}$ is the embedding landscape calculated on category $c$ according to Equation 1. Given the neuron contribution set $\mathbf{C}^{l,c}=\{\mathbf{C}_{n_{0}}^{l,c},\mathbf{C}_{n_{1}}^{l,c},...,\mathbf{C}_{n_{|\mathbf{N}^{l}|-1}}^{l,c}\}$, we can further reformulate the given problem into a variant of the Prefix Sum Problem 222 The Prefix Sum problem, in its simplest form, involves computing an array of cumulative sums (prefix sums) of a given sequence of numbers.. Specifically, the objective is to identify a particular value of $k$ such that $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$ given the layer-wise contribution value set $\mathbf{C}^{{l,c}}$.
Algorithm Implementation. To address the aforementioned issue, we adopt a Linear Scan strategy as described by Algorithm 1. Specifically, this approach starts with the neurons possessing the top $i$ highest contribution scores in the array $\mathbf{C}^{l,c}$. We initially set $\mathbf{\Omega}^{l,c}$ to encompass these top $i$ neurons and proceed to calculate the similarity score. For each neuron contribution score within $\mathbf{C}^{l,c}$, the strategy involves sequentially selecting the top $i$ neurons with the highest contribution scores from the neurons not in $\mathbf{\Omega}^{l}$ (i.e., $\mathbf{N}^{l}\setminus\mathbf{\Omega}^{l,c}$). This process is continued until the condition $\text{sim}_{\text{sem}}(f^{\mathbf{\Omega}_{k}^{l,c}},f^{l,c})>\Theta$ is met.
To reduce the computational burden involved in incrementally adding neurons to the subset and computing their semantic similarity with all neurons in the layer, we can increase the value of $i$.
Data: All neurons $\mathbf{N}^{l}$ of layer $l$, the neuron contribution set $\mathbf{C}^{l,c}$, the semantic similarity threshold $\Theta$, the selection interval $i$.
Result: The critical neuron set for category $c$: $\mathbf{\Omega}^{l,c}$.
1 $\Omega^{l,c}\leftarrow\emptyset$
2while $\mathrm{sim}_{\mathrm{sem}}(f^{\Omega^{l,c}},f^{l,c})<\Theta$ do
3       $\Omega^{l,c}\leftarrow\Omega^{l,c}\cup\text{top}i(\mathbf{N}^{l}\setminus\Omega^{l,c},\mathbf{C}^{l,c})$
Algorithm 1 Linear Scan Algorithm.
Categorical Critical Neurons Aggregation. After determining the critical neurons for each category, they can be combined on a category basis to establish the critical neuron set for the entire model, as shown in Figure 1. For an entire dataset $D$ containing $m$ categories, this can be expressed as: $\mathbf{\Omega}^{l,D}=\mathbf{\Omega}^{l,D_{c_{0}}}\cup\mathbf{\Omega}^{l,D_{c_{1}}}\cup\ldots\cup\mathbf{\Omega}^{l,D_{c_{m-1}}}$. We can further depict the neuron contribution of the model $f$ as $\mathbf{\Omega}=\{\mathbf{\Omega}^{l_{0,D}},\mathbf{\Omega}^{l_{1,D}},...,\mathbf{\Omega}^{l_{L-1,D}}\}$. Given the neuron contribution metrics $\mathbf{C}$, the selection threshold $\Theta$ directly determines $\mathbf{\Omega}$, simplifying the process by eliminating the need to consider varying optimal neuron proportions across categories and layers.
4. semantic-aware Model Maintenance
4.1. Model Restructure
Identifying critical neurons enables the removal of non-critical ones, keeping only essential neurons in each layer. This reduces the model’s size by eliminating ‘redundant’ elements while preserving prediction accuracy, as shown in Figure 2. This process uses an indicator function $I(n)$ for each neuron $n$ to decide whether to retain it based on its contribution. The model restructure strategy can be expressed as:
$I(n)=\left\{\begin{array}[]{ll}1&\text{if }n\in\mathbf{\Omega},\\ 0&\text{otherwise}.\end{array}\right.$
Then each neuron $n^{\prime}$ in the restructured model $f^{\prime}$ could be expressed as $n^{\prime}=n\cdot I(n)$. Note that the threshold value $\Theta$ can be dynamically adjusted to balance model performance and size during restructuring. By adjusting $\Theta$, this method allows for a restructured model that varies in accuracy and size to suit diverse needs. Such restructuring will lead to an effective decrease in the storage space. Moreover, some weight-level unstructured pruning techniques, such as (Lee et al., 2020; Morcos et al., 2019; Gale et al., 2019; Evci et al., 2019), can be combined with our method. By leveraging model restructure, these pruning methods can focus on redundant weights within critical neurons, thus accelerating the pruning process and further reducing the model size.
Figure 2. Model Restructure based on semantic slicing. We only retain the critical neurons for model restructure.
4.2. Model Re-adaption
Model re-adaptation improves adaptability and robustness by focusing on training parameters related to critical neurons. Previous research (Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021) has shown that in cases of poor performance, such as adversarial examples, adversarial noise is amplified and propagated through critical neurons. Therefore, the fundamental cause is the abnormal behavior of critical neurons leading to errors, while other neurons are less affected. Retraining these critical neurons’ parameters can thus more efficiently repair and enhance the model in terms of cases with poor performance. The complete model re-adaption workflow is shown in Figure 3. Using adversarial samples as an example, the retraining strategy adopts the following loss function:
$\mathcal{L}=\mathcal{L}_{c}(x,y;\theta)+\mathcal{L}_{c}(adv_{x},y;\theta),$
where $L_{c}$ denotes the cross entropy, $adv_{x}$ is the adversarial counterpart of the sample $x$, and the notation $\theta$ represents the parameters in model $f$. The parameters update strategy in the retraining is:
$\theta^{\prime}_{\mathbf{\Omega}}=\theta_{\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{\Omega}}}\mathcal{L},$
where $\eta$ is the learning rate. Note that in this strategy, the parameters that fall outside the critical neurons set $\mathbf{\Omega}$ remain unchanged and are not updated.
Figure 3. Model Re-adaption based on semantic slicing. We train only the critical neurons for model re-adaptation.
4.3. Model Incremental Development
The model incremental development refers to the DNN model to learn from new data over time without forgetting previously learned information. This is crucial for creating DNN models that can adapt and evolve in dynamic environments. Taking an image classification designed to recognize different types of animals as an example, we initially only have data to train a DNN to classify images into three categories: Dogs, Cats, and Birds (old task). After the initial training, we collect more data and want the DNN to learn to recognize two new categories Fish and Horses (new task). It is a significant challenge for DNN models. ❶ In practical applications, the training data for the old task might be discarded to save storage space and we only have access to the data of the new task. When training only with the data of the new task, the incremental development of DNNs is prone to ‘catastrophic forgetting’, where learning new tasks can lead to the loss of old tasks. ❷ Even with the training data of the old task, the training process is less resource-efficient because the DNN models have to be retrained with the entire dataset (new and old data).
This issue is particularly prominent in real-world scenarios, such as in IoT devices (Liu et al., 2021), where incremental development is essential. Due to space constraints, these devices might initially be trained on a limited set of data categories. However, as new data becomes available, it is often impractical to retrain the entire model from scratch or to delete previously learned categories.
Our strategy achieves model incremental development by freezing critical neurons’ weights during new task learning, preserving performance on original tasks while allowing new skills acquisition by training the model’s ‘redundant’ components. Specifically, for a model that has already been trained for task $A$, our initial step involves identifying the critical neurons associated with task $A$. Subsequently, when training for a new task, task $B$, we fix the parameters related to these critical neurons as shown in Figure 4. This approach allows the model to learn $B$ by focusing on training the non-critical neurons. Since the critical neurons and their parameter settings preserve the critical information of task $A$, we maintain the predictive accuracy for task $A$ by retaining the critical neurons.
The model incremental development strategy follows the following loss:
$\mathcal{L}=\mathcal{L}_{c}(x_{B},y_{B};\theta),$
where $L_{c}$ denotes the cross entropy and $(x_{B},y_{B})$ represents data pairs from task B. The parameters update strategy is:
$\theta^{\prime}_{\mathbf{N}\setminus\mathbf{\Omega}}=\theta_{\mathbf{N}\setminus\mathbf{\Omega}}-\eta\cdot\nabla_{\theta_{\mathbf{N}\setminus\mathbf{\Omega}}}\mathcal{L},$
where $\eta$ is the learning rate, $\mathbf{N}$ represents all neurons in model $f$, and $\mathbf{N}\setminus\mathbf{\Omega}$ represents the neurons outside the critical neurons set $\mathbf{\Omega}$. In the inference process, we will exclusively utilize the frozen parameters associated with task $A$ for inference on task A, and employ the remaining parameters for task B.
Figure 4. Model Incremental Development based on semantic slicing. We train only the non-critical neurons for model incremental development.
5. Evaluation
In this section, we evaluate the performance of NeuSemSlice. We first outline the experimental setup and then, we introduce our evaluation aiming to answer the following research questions:
  • RQ1: How does semantic slicing outperform other DNN slicing methods?
  • RQ2: Is NeuSemSlice effective for model restructuring?
  • RQ3: Is NeuSemSlice effective for model re-adaption?
  • RQ4: Is NeuSemSlice effective for model incremental development?
5.1. Evaluation Setup
To ensure the comprehensiveness and depth of the experiments, we select four datasets widely utilized in the deep neural network domain, along with five network models for experimental evaluation. ❶ MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017). For training, we use a Multilayer Perceptron (MLP) model with two hidden layers, containing 512 and 256 neurons respectively. ❷ CIFAR-10 (Krizhevsky et al., 2009). Three deep neural network models—AlexNet (Krizhevsky et al., 2012), VGG-16 (Simonyan and Zisserman, 2014), and ResNet-101 (He et al., 2016)—are utilized for training. ❸ ImageNet (Deng et al., 2009). Following the setup in (Xie et al., 2022), only data from the top 10 categories are selected for experiments, with the VGG-19 model employed.
Based on the datasets and models mentioned, six experimental settings are constructed: MLP-MNIST, MLP-FMNIST, AlexNet-CIFAR10, VGG16-CIFAR10, ResNet101-CIFAR10, and VGG19-ImageNet. These experiments include a variety of configurations, ranging from small to large models and datasets. In each experimental setting, we train at least five different models to repeat the experiments to reduce randomness and ensure robustness. The results are presented as mean values to enhance the experiment reliability and precision.
All these experiments are conducted on the Intel Xeon Silver 4214 Processor with 2 Tesla V100 GPUs with 32GB memory. We have implemented our tool based on Pytorch (Paszke et al., 2019a). The source code, related information of datasets and models, and complete experimental results are released at (NeuSemSlice, 2024).
5.2. RQ1: The performance of semantic slicing
In RQ1, we aim to evaluate and compare various slicing methods. Specifically, we follow (Xie et al., 2022; Li et al., 2021) to mask the redundant components of the DNN model (i.e., zeroing non-critical neurons) and then compare the model accuracy with only the critical neurons retained. Greater accuracy suggests the slicing method could better identify critical neurons with precision.
Figure 5. Comparison results of performance across different slicing methods are presented. Here, we only present the results for the top 3 contribution metrics for better observation. Each dot in the figure represents the model performance (i.e., compression rate and accuracy) after masking the redundant components of neurons identified by a slicing method, with DeepArc represented by a ★. The lines representing semantic slicing are plotted by varying $\Theta$ values from 0.9 to 0.99. For global and category-wise slicing, the top 10%, 20%, up to 100% of neurons per layer or category are selected for slicing respectively. The results for DeepArc are obtained by removing redundant layers with semantic similarity greater than 0.99. Ideally, dots (lines) located at the upper right are favored for their higher compression rates and better accuracy performance, which also suggest that the corresponding slicing methods better identify critical neurons in the model.
Detailed Configurations. In RQ1, we compare the following three slicing methods: global slicing, category-wise slicing, and semantic slicing. To assess neuron contribution and select critical neurons, we utilize five neuron contribution metrics: Avg, Var, Ens, DeepLIFT, and Taylor. As outlined in Section 3.1, global slicing and category-wise slicing choose top 10%, 20%, … up to 100% of neurons in each layer or category for slicing respectively. Additionally, we also introduce DeepArc for comparative analysis. Following the settings used in (Ren et al., 2023), we set the threshold at 0.99 to identify redundant layers in the model. Subsequently, we remove these redundant layers and compare the model accuracy when only critical semantic layers are retained.
Semantic slicing identifies critical neurons for each category using a set of thresholds $\Theta=0.9,0.91,0.92,0.93,0.94,0.95,0.96,0.97,0.98,0.99$, following the similar settings used by (Ren et al., 2023). Then it merges them into semantic slices, as described in Section 3.3. For RQ1, the compression rate is defined as the percentage reduction of parameters in the model, while model accuracy denotes the predictive accuracy of the model slices on the test dataset. We conducted comparisons of slicing precision across the previously mentioned six experimental settings.
Comparison with Global Slicing and Category-wise Slicing. Figure 5 compares the slicing precision differences among semantic slicing, global slicing, and category-wise slicing (The complete experimental results can be found in (NeuSemSlice, 2024)). For complex models like AlexNet and VGG16, global slicing and category-wise slicing can only maintain certain model performance at very low compression rates, with accuracy significantly declining as the compression rate increases. Semantic slicing demonstrates significant improvement over the other two techniques across all five neuron contribution metrics. For example, in the AlexNet-CIFAR10 case (Figure 5(c)), semantic slicing uses the neuron contribution metric Taylor to maintain model performance, with only a 0.38% loss in accuracy at a 50% compression rate. Conversely, the other two techniques experience significant accuracy losses of around 70% at the same compression rate. Similarly, on VGG16 using the DeepLIFT metric (Figure 5(d)), semantic slicing at around a 10% compression rate reduces accuracy loss by 41.18% and 77.77% compared to global slicing and category-wise slicing.
On larger models like ResNet101 and VGG19 (Figure 5(e) and (f)), global slicing and category-wise slicing are also less effective. Even for datasets like MNIST and Fashion-MNIST (Figure 5(a) and (b), which require fewer neurons for high accuracy, semantic slicing still slightly outperforms them, maintaining higher model accuracy across all evaluated neuron contribution metrics at the same levels of compression rate.
These results adequately confirm that inadequate account for the uneven distribution of critical neurons across different categories and network layers would lead to poor slicing performance. This highlights the advantage of the proposed technique, which can dynamically identify critical features and components of the model based on semantic similarity.
The Effects of DeepArc. For DeepArc, we identify redundant layers in six settings based on layer similarity, as shown in Figure 5. Firstly, for the MLP, the model only has two hidden layers. When any layer is removed, we follow the model restructuring detailed in DeepArc and introduce a “sewing layer” to align the input and output of the layers. However, this results in no change in the size of the model, indicating that DeepArc cannot effectively identify redundancies in such models. In the other four configurations, DeepArc only identified redundant layers with a similarity greater than 0.99 in VGG16 and ResNet101.
For AlexNet, the maximum layer similarity is 0.95. If the threshold is lowered to 0.95 and the corresponding redundant layer in the module is removed, a fully connected layer containing 87.97% of the model’s parameters is eliminated, reducing the model accuracy to 10.94%. The results for VGG19 are similar, with the highest layer similarity being only 0.97. For VGG16, removing redundant layers within the module leads to a model compression rate of 32.06% but an accuracy of only 14.84%. For ResNet101, after removing 10.68% of redundant layers, the accuracy remains at 84.77%. However, its precision in identifying critical model components is still not as high as that of NeuSemSlice. These results suggest that DeepArc is relatively effective only in large models with small datasets, due to the difficulty in sharing semantics in smaller models or models with lower redundancy.
Comparing different Neuron Contribution Metrics. Our NeuSemSlice framework is orthogonal to the neuron contribution metrics, meaning it can be integrated seamlessly with a range of contribution metrics. Through our framework, we can fairly compare different neuron contribution metrics. We can observe that in relatively simple models like MLP, the performance of these metrics is close. They all effectively assess neuron contributions, enabling the precise selection of critical neurons. For example, in the case of MLP-MNIST, even at a compression rate of 70%, the model’s accuracy only drops by approximately 4%. In contrast, for more complex models such as AlexNet and VGG16, DeepLIFT and Taylor metrics perform better. For AlexNet, the Taylor metric achieves a mere 3% accuracy drop at a compression rate of 66.05%; for VGG16, both DeepLIFT and Taylor metrics maintain the model’s accuracy within a 2% decrease at a compression rate of approximately 10%. For larger models like ResNet101 and VGG19, the performance of the metrics is similarly close in ResNet101, with Ens achieving only a 1.54% drop in accuracy at a 10.56% compression rate, while in VGG19, Taylor significantly outperforms other metrics, maintaining a 2.54% accuracy drop at a 24.67% compression rate.
The Var metric exhibits lower accuracy after model compression, which is especially evident in AlexNet and VGG16. We also calculate the time required to compute various metrics: Avg, Var, Ens, DeepLIFT, and Taylor, which take 5.4s, 5.5s, 5.6s, 30.9s, and 6.2s respectively on the MNIST dataset. In comparison to other methods, DeepLIFT exhibits significantly higher time consumption. Overall, the DeepLIFT and Taylor metrics outperform the other three metrics, while DeepLIFT is less efficient.
Answer to RQ1: Compared to global slicing and category-wise slicing, semantic slicing better identifies critical neurons with precision. At 50% compression rate, semantic slicing increases model performance from 23.24% and 11.55% to 89.20% compared to two other techniques respectively (see AlexNet-CIFAR10, neuron contribution metric Taylor, Figure 5(c)).
5.3. RQ2: Model Restructuring
Table 4. Results of Model Restructuring.
MLP-MNIST MLP-FMNIST AlexNet-CIFAR10 VGG16-CIFAR10 ResNet101-CIFAR10 VGG19-ImageNet
Compression
Algorithm
CR $\uparrow$ NI $\downarrow$ MA $\uparrow$ CR $\uparrow$ NI $\downarrow$ MA $\uparrow$ CR $\uparrow$ NI $\downarrow$ MA $\uparrow$ CR $\uparrow$ NI $\downarrow$ MA $\uparrow$ CR $\uparrow$ NI $\downarrow$ MA $\uparrow$ CR $\uparrow$ NI $\downarrow$ MA $\uparrow$
LAMP 52.17% 7 98.26% 36.97% 9 90.01% 30.17% 7 89.49% 18.55% 4 91.28% 45.96% 12 90.77% 30.17% 7 89.54%
NeuSemSlice+LAMP 53.05% 5 98.32% 39.70% 8 90.21% 34.88% 6 89.52% 20.36% 4 91.29% 46.46% 10 92.11% 37.25% 6 89.19%
Glob 40.95% 5 98.31% 33.66% 8 90.09% 60.28% 18 89.44% 22.62% 5 91.25% 53.67% 15 72.21% 22.62% 5 89.42%
NeuSemSlice+Glob 42.04% 3 98.37% 36.52% 7 90.18% 62.96% 17 89.44% 24.34% 5 91.28% 54.10% 13 73.96% 26.93% 3 89.12%
Unif 40.95% 5 98.34% 40.13% 10 89.81% 55.99% 16 88.59% 14.26% 3 91.24% 51.23% 14 83.95% 55.99% 16 85.85%
NeuSemSlice+Unif 42.04% 3 98.36% 42.71% 9 89.95% 58.96% 15 88.68% 16.17% 3 91.31% 51.68% 12 84.03% 56.25% 13 87.50%
ERK 34.39% 4 98.30% 30.17% 7 90.08% 55.99% 16 89.37% 26.49% 6 91.19% 14.26% 3 95.66% 26.49% 6 89.58%
NeuSemSlice+ERK 42.04% 3 98.33% 33.18% 6 90.24% 58.96% 15 89.41% 28.13% 6 91.28% 23.34% 3 93.97% 34.06% 5 89.19%
Detailed Configuration. To systematically assess NeuSemSlice for the model restructuring task, we adopt the same experimental setup detailed in (Ren et al., 2023), and validate by comparing the compression efficiency and effects on the model before and after applying NeuSemSlice. Specifically, after removing the non-critical neurons of the model, we expect the compression technology to directly focus on the redundant parts within the critical neurons, improving the iteration efficiency [^3] It is important to note that in this experiment and the subsequent RQ3 and RQ4, our method is based on the optimal configuration identified in RQ1. The term “our method combined with DeepArc” refers to retraining the critical neurons identified by our method in the critical modules recognized by DeepArc (targeting adversarial samples).. Based on this, we apply four of the most advanced model compression techniques currently available—LAMP (Lee et al., 2020), Global (Morcos et al., 2019), Uniform+ (Gale et al., 2019), and ERK (Evci et al., 2019)—for optimal compression results.
In our experiments, we take the following metrics to evaluate the performance of model compression. The Compression Rate (CR) indicates the reduction in parameter size, the Number of Iterations (NI) reflects the efficiency of the compression process, and Model Accuracy (MA) measures the predictive performance on the test dataset.
Results and Findings. Table 4 presents the experimental results of NeuSemSlice in model restructuring. Across all experimental settings, the application of NeuSemSlice in model restructuring allows models to achieve higher compression ratios with fewer iterations, indicating that NeuSemSlice can improve both compression efficiency and effectiveness. Apart from the VGG19-ImageNet configuration, NeuSemSlice not only enhances compression efficiency and rates but also slightly improves model accuracy. Specifically, in six experimental settings, the average number of iterations required for model compression decreases from 8.67 to 7.50, enhancing efficiency by 13.46%, and the model compression ratio increases from 37.26% to 40.26%. These results demonstrate that the application of NeuSemSlice in model restructuring boosts the performance of model compression and enhances the performance of compressed models.
Answer to RQ2: For model restructuring, NeuSemSlice not only elevates the efficiency of compression technology—an average increase of 13.46%—but also achieves approximately a 3% enhancement in the compression rate and a slight increase in accuracy.
5.4. RQ3: Model Re-adaptation
Table 5. Results of Model Readaptation.
MLP-MNIST MLP-FMNIST AlexNet-CIFAR10 VGG16-CIFAR10 ResNet101-CIFAR10 VGG19-ImageNet
Method TPR $\downarrow$ RA $\uparrow$ RAA TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$ TPR $\downarrow$ RA $\uparrow$ RAA $\uparrow$
FGSM Fully Training 100.00% 100.00% 14.20% 100.00% 98.25% 10.10% 100.00% 98.40% 43.90% 100.00% 97.83% 13.10% 100.00% 100.00% 49.50% 100.00% 98.54% 48.59%
DeepArc 75.01% 100.00% 13.40% 75.01% 97.18% 11.20% 99.18% 99.43% 50.40% 67.60% 98.80% 13.70% 56.26% 100.00% 53.60% 95.82% 98.44% 42.66%
NeuSemSlice 26.74% 100.00% 18.10% 49.26% 97.72% 11.40% 33.95% 99.55% 46.10% 87.41% 99.60% 15.70% 86.35% 100.00% 54.20% 85.22% 98.78% 47.66%
NeuSemSlice+DeepArc 24.17% 100.00% 21.20% 40.29% 98.00% 11.50% 33.21% 100.00% 52.70% 60.39% 99.88% 37.30% 47.30% 100.00% 54.60% 81.12% 98.85% 44.22%
PGD Fully Training 100.00% 100.00% 14.00% 100.00% 97.90% 6.00% 100.00% 97.68% 1.30% 100.00% 95.80% 2.90% 100.00% 99.00% 0.90% 100.00% 99.38% 22.81%
DeepArc 75.01% 100.00% 12.00% 75.01% 97.37% 7.20% 99.18% 99.95% 0.10% 67.60% 99.53% 0.00% 56.26% 95.30% 4.20% 95.82% 99.11% 20.78%
NeuSemSlice 26.74% 100.00% 12.90% 49.26% 97.90% 5.80% 33.95% 99.30% 1.40% 87.41% 99.75% 0.00% 86.35% 99.40% 1.40% 85.22% 99.18% 20.16%
NeuSemSlice+DeepArc 24.17% 100.00% 12.60% 40.29% 97.92% 8.10% 33.21% 98.98% 0.30% 60.39% 99.83% 0.00% 47.30% 98.93% 4.40% 81.12% 99.41% 24.06%
Detailed Configuration. To answer RQ3, we conduct an experiment using two prevalently-employed adversarial sample generation techniques, Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) and Projected Gradient Descent (PGD) (Madry et al., 2017). We generate 1,000 adversarial samples, with FGSM reducing the model’s average accuracy to 18.53%, and PGD to 3.8%. We aim to compare the Training Parameter Ratio (TPR) and the Retrained Accuracy (RA) of existing approaches and our proposed method. Specifically, the experiment settings include fully training (i.e., comparison baseline), DeepArc, NeuSemSlice, and a combination of NeuSemSlice with DeepArc under identical training epochs. Additionally, this RQ further assesses the resilience of each method against similar adversarial attacks. We generate another 1,000 adversarial samples after the model re-adaption, and obtain the model accuracy against the samples (i.e., shown as Re-Attack Accuracy, RAA).
Results and Findings. Table 5 presents our experimental results in the four settings. Compared to the traditional method of fully training) and DeepArc, our method demonstrates a significant reduction in the required model parameters for training, averaging only 61.49%. This represents an average decrease of 38.51% in training parameters compared to the fully training strategy and an average reduction of 16.66% reduction relative to DeepArc. Moreover, our method surpasses fully training by an average of 0.67% and DeepArc by an average of 0.50% in Retrained Accuracy, demonstrating advantage in both training cost-efficiency and effectiveness. Among the four methods, our method combined with DeepArc achieves a 0.73% increase in Retrained Accuracy compared to traditional training methods, with the lowest training cost (averaging only 47.75% of the parameters trained).
Our method combined with DeepArc also shows the best performance in terms of re-attack performance, achieving up to a 17.8% improvement in RAA. The results further demonstrate the superior performance of our method in model re-adaptation scenarios. However, we observe that all of the methods experience notable performance loss against re-attack, alerting us to design better methods against adversarial attacks in the future, to be discussed in Section 5.6. Finally, NeuSemSlice is comparable to DeepArc in efficiency, requiring only 93.53% of the time needed for full training on average. When NeuSemSlice is combined with DeepArc, it requires only 89.80% of the full training time. The detailed data is available at (NeuSemSlice, 2024).
Answer to RQ3: Our approach reduces the cost of model re-adaptation by maintaining critical neurons, requiring only an average of 61.49% of the parameters. Our method also exhibits slight advantages in terms of efficiency and model accuracy over baselines.
5.5. RQ4: Model Incremental Development
Table 6. Results of Model Incremental Development.
MLP-MNIST MLP-FMNIST AlexNet-CIFAR10 VGG16-CIFAR10 ResNet101-CIFAR10 VGG19-ImageNet
Method TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$ TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$ TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$ TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$ TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$ TaskA $\uparrow$ TaskB $\uparrow$ Avg $\uparrow$
Original 98.92% 0.00% 49.46% 97.44% 0.00% 48.72% 89.14% 0.00% 44.57% 92.52% 0.00% 46.26% 96.90% 0.00% 48.45% 92.15% 0.00% 46.08%
Retrain 0.00% 98.50% 49.25% 2.78% 83.28% 43.03% 0.00% 89.40% 44.70% 0.00% 90.80% 45.40% 0.00% 85.86% 42.93% 5.37% 88.92% 47.15%
Replay 5% 87.54% 98.32% 92.93% 83.96% 83.14% 83.55% 46.32% 88.94% 67.63% 54.42% 90.76% 72.59% 26.54% 85.00% 55.77% 85.62% 87.77% 86.69%
Replay 10% 91.86% 98.13% 95.00% 87.90% 83.20% 85.55% 59.70% 88.16% 73.93% 62.80% 90.12% 76.46% 39.40% 84.68% 62.04% 87.23% 88.38% 87.81%
Replay 100% 97.20% 97.48% 97.34% 94.44% 81.78% 88.11% 82.96% 85.66% 84.31% 86.98% 88.30% 87.64% 84.12% 90.18% 87.15% 90.46% 89.38% 89.92%
DeepArc / / / / / / 0.00% 84.84% 42.42% 20.44% 0.00% 10.22% 46.68% 0.00% 23.34% 0.00% 84.92% 42.46%
NeuSemSlice 98.89% 97.20% 98.04% 92.12% 82.08% 87.10% 84.65% 87.60% 86.13% 85.77% 86.98% 86.37% 97.12% 82.52% 89.82% 91.77% 90.54% 91.15%
NeuSemSlice+Replay 10% 98.94% 97.17% 98.05% 96.96% 81.90% 89.43% 87.22% 86.66% 86.94% 91.38% 86.02% 88.70% 98.52% 81.46% 89.99% 92.54% 90.38% 91.46%
Detailed Configuration. To answer RQ4, we split the datasets MNIST, FMNIST, CIFAR10, and ImageNet into two subtasks, each with five categories, labeled as task $A$ and task $B$. For instance, in the CIFAR10 dataset, which contains ten distinct image categories, task $A$ consists of the first five categories, while task $B$ consists of the remaining five categories.
To simulate the process of incremental model development, we start by training models exclusively on task $A$ in six different experimental setups following the division. Then, we proceed to train these models on task $B$ while aiming to preserve their performance on task $A$. The evaluation focuses on the accuracy for both task $A$ and task $B$. Here, the accuracy on task $A$ evaluates how well the models retain knowledge of the original task without forgetting, while the accuracy on task $B$ gauges the models’ ability to learn new tasks.
We would like to note that this setup aims to simulate the actual scenario of model incremental development, where we typically have limited access to the training data of task $A$. Therefore, we introduce ‘replay’ as a baseline for comparison, which refers to relaxing the restrictions on accessing the data of task $A$ while learning task $B$. This is done by revisiting some of the task $A$ data to aid the model in remembering and maintaining its performance on the old task. In particular, we aim to compare the following seven methods: naive retraining (i.e., directly training with the data of task B), Replay 5% (i.e., the training data includes 5% of task $A$ data with all data of task $B$), Replay 10%, Replay 100%, DeepArc (i.e., training task $B$ only on the redundant layers of the model), NeuSemSlice, and NeuSemSlice combined with Replay.
Results and Findings. The results, as shown in Table 6, initially reveal that direct retraining causes catastrophic forgetting in every setting, with the model completely forgetting knowledge related to task $A$ in four settings. Replay methods outperform direct retraining, with 5% data replay resulting in an average 30.45% decrease in predictive accuracy for task $A$, 10% data replay leading to a 23.03% average decrease, and 100% data replay causing a 5.15% average decrease.
Our method significantly outperforms the 5% and 10% data replay and even surpasses the 100% data replay in four experimental settings. The average precision loss for task $A$ is only 2.79%, with an average accuracy advantage of 0.69% over the 100% data replay. When combined with 10% data replay, our method consistently achieves the best results in all experimental settings, with an average precision loss for task $A$ of only 0.25% and an average accuracy of 90.76%.
These results thoroughly demonstrate that our method effectively mitigates the issue of catastrophic forgetting in task $A$. Even with a fixed portion of the model’s parameters, our method incurs only around a 1.64% loss in accuracy on task $B$ compared to direct retraining. Notably, in terms of storage efficiency, replay methods require storing some training data from task $A$, typically needing several to tens of megabytes of storage space under our experimental settings. In contrast, our method only needs to retain index information of critical neurons in the model, significantly reducing storage requirements to just a few kilobytes, which is particularly crucial for applications with limited storage resources. Moreover, in terms of efficiency, our method is slightly superior to the 5% and 10% data replay and is twice as efficient as the 100% data replay, with more details provided in (NeuSemSlice, 2024). Overall, considering the combined performance on tasks $A$ and $B$, our method demonstrates significant effectiveness in the model incremental development task.
Regarding DeepArc, we find that training on the identified redundant layers of the model does not effectively alleviate catastrophic forgetting. It maintains only partial capability to complete task $A$ on VGG16 and ResNet101, and DeepArc leads to a substantial reduction in the model’s ability to learn new tasks. A possible explanation is that training only the model’s redundant layers is insufficient for learning new tasks.
Answer to RQ4: Our method reduces the average accuracy loss for the original task to 2.79% and achieves 89.77% average accuracy in the incremental model development task. Combined with replay, it consistently excels across all settings, reducing the average accuracy loss for the original task to 0.25% and increasing average accuracy to 90.76%, effectively countering catastrophic forgetting.
5.6. Threats to Validity
5.6.1. External Threats
The threat to external validity arises from the particular settings we have chosen for our study, which means raising the generalizability concern of our proposed approach. To mitigate these threats, we have selected diverse evaluation settings. In more detail, we select four widely utilized datasets–MNIST, Fashion-MNIST, CIFAR10, and ImageNet, and employ five classical DNN architectures– MLP, AlexNet, VGG16, ResNet101, and VGG19. The experiments are conducted under six settings. This selection is intended to enhance the generalizability of our findings across common scenarios within the domain.
5.6.2. Internal Threats
Internal validity threats stem from the tools utilized in our research, which include captum (Kokhlikyan et al., 2020) and cleverhans (Papernot et al., 2018). Additionally, there is a potential threat concerning the accurate reproduction of the contribution assessment metrics and the model modularization framework DeepArc in PyTorch (Paszke et al., 2019b). The use of various deep learning frameworks, such as TensorFlow (Abadi et al., 2015), may impact the results. Additionally, the inherent randomness in model training poses a potential threat to the internal validity. We addressed this issue by repeating key experiments more than five times and reporting the mean values. Finally, as mentioned in Section 5.4, the resilience of the model’s re-adaptation to adversarial attacks requires enhancement. This issue, though orthogonal to the main focus of this paper, highlights the need for future work to tackle and reduce its impact.
6. Conclusion
This paper introduces NeuSemSlice, an innovative framework that integrates semantic slicing with semantic-aware model maintenance for deep neural networks. Semantic slicing identifies critical neurons at a granular level, categorizing and merging them across layers based on semantic similarities. This approach markedly enhances the effectiveness of these neurons in model maintenance tasks, surpassing global slicing and category-wise slicing. NeuSemSlice further innovates in semantic-aware model maintenance by offering strategies for model restructuring that preserve critical neurons, efficient model readaptation through tuning only these neurons, and incremental model development that maintains critical neurons while training non-critical ones. The effectiveness of NeuSemSlice is demonstrated through comprehensive evaluations, showing its superiority over baselines in all three aspects of model maintenance.